OpenStudy (anonymous):

If the amount of air in a spherical balloon is 288 cu. in., what is the radius of the balloon?

OpenStudy (lgbasallote):

Volume of a sphere is \[\huge V = \frac 43 \pi r^3\] so to find r just do \[\huge r = \sqrt[3]{\frac{3V}{4\pi}}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yes , thank you !